OpenStudy (anonymous):

1. The equation of a circle is shown below. (x -5)2 + (y + 2)2 = 64 The radius of the circle is ________ units. (1 point)

OpenStudy (gorv):

8

OpenStudy (gorv):

(x-x1)^2+(y-y1)^2=r^2

OpenStudy (gorv):

@animatedstar

OpenStudy (anonymous):

so is the answer 8 units

OpenStudy (anonymous):

@animatedstar the general equation of a circle is (x-a)^2 + (y-b)^2 = r^2 where (a,b) are the coordinates of the centre and r is the radius of circle

OpenStudy (gorv):

yes

OpenStudy (anonymous):

does anyone know the know the answer

OpenStudy (gorv):

8 is correct

OpenStudy (anonymous):

thx can i ask you another question

OpenStudy (anonymous):

What is the measure of angle F? (1 point)

OpenStudy (gorv):

put sum of angles =720

OpenStudy (gorv):

if u got plzz click best response

OpenStudy (anonymous):

so the measure of angle f is 720

OpenStudy (gorv):

30x+8+134+127=720 30x=720-169 x=18.36

OpenStudy (gorv):

f=10x-4=183.6-4=179.6

OpenStudy (anonymous):

so its 179.6

OpenStudy (gorv):

yes